# CAVE: Controllable Authorship Verification Explanations

### v2

@@ Abstract
Authorship Verification (Av) (do two documents have the same author?) is essential in many sensitive real-life applications. Av is often used in proprietary domains that require a private, offline model, making SOTA online models like ChatGPT undesirable. Current offline models however have lower downstream utility due to low accuracy/scalability (eg: traditional stylometry Av systems) and lack of accessible post-hoc explanations. In this work, we take the first step to address the above challenges with our trained, offline LLaMa-3-8B model Cave (Controllable Authorship Verification Explanations): Cave generates free-text Av explanations that are controlled to be (1) structured (can be decomposed into sub-explanations in terms of relevant linguistic features), and (2) easily verified for explanation-label consistency (via intermediate labels in sub-explanations). We first engineer a prompt that can generate silver training data from a SOTA teacher model in the desired Cave output format. We then filter and distill this data into a pretrained LLaMa-3-8B, our carefully selected student model. Results on three difficult Av datasets IMDb62, Blog-Auth, and FanFiction show that Cave generates high quality explanations (as measured by automatic and human evaluation) as well as competitive task accuracies111github.com/INK-USC/Controllable-AV-Explanations .
@@
Sahana $\varheartsuit\varheartsuit{}^{\varheartsuit}$, Kartik $\varheartsuit\varheartsuit{}^{\varheartsuit}$, Elizabeth Boschee, Xiang $\varheartsuit\varheartsuit{}^{\varheartsuit}$$\varheartsuit{}^{\varheartsuit}$epartment of Computer Science, University of Southern California Information Sciences Institute, University of Southern California sramnath@usc.edu
@@ 1 Introduction
Figure 1: Cave generates structured free-text explanations that can be automatically verified for consistency.
@@
Authorship Verification (Av) Koppel and Winter (2014) is the NLP task of determining if two input documents were written by the same author. Av is used for tasks such as plagiarism detection, forensic analysis (often in support of law enforcement), analysis of the spread of misinformation. Given the sensitive nature of this task, it is imperative to develop Av methods that are explainable, accessible and secure. Explainability ensures that the system’s decisions can be trusted by user. Accessibility ensures easier downstream usability; for example, a well-structured free-text explanation with clearly defined topic points is more easily understood than unstructured explanations such as long lists of low-level features. Finally, security is necessary to safely use the Av system for proprietary data; a locally-hosted model on the user’s server is more secure than an API call to an online model, which could potentially store the user’s sensitive data. Av is also an extremely difficult task for humans, in contrast to other NLP tasks such as machine translation; this further necessitates the need for automatic Av systems that accessible to the user, accurate and explainable.
@@
Early works performed Av by comparing hand-crafted features such as n-grams van Dam (2013), POS tags Moreau and Vogel (2013), LIWC features Uchendu et al. (2020); while these are explainable, they are hard to scale. Subsequent works employed neural architectures to improve scalability and accuracy, such as the usage of Siamese networks Araujo-Pino et al. (2020); Najafi and Tavan (2022); Boenninghoff et al. (2019) to assign authorship via distance between document embeddings. These methods, however, have little or no post-hoc explainability. This poses a huge problem: one cannot, for instance, make legal assertions about the authorship of a set of questionable documents solely on the basis of “a system said so” - decision makers (judges, university officials, intelligence analysts, etc.) need to know why.
Recent years have seen the surge of large language models and their self-rationalization capabilities. Subsequently, there have been works Hung et al. (2023); Huang et al. (2024) that generate free-text rationales [^2] We use terms explanation and rationale interchangeably. for Av from models such as GPT-4 OpenAI (2024b) in a zero/few-shot manner. These methods however face issues such as unreliability of online models (varying results/details across runs, deprecation of APIs), inconsistency of rationales with predicted label Hung et al. (2023) (refer to Appendix B), and high expenses.
@@
In this work, we propose our model Cave (Controllable Authorship Verification Explanations), an in-house model that generates authorship predictions along with high-quality, free-text rationales that explain the former post-hoc. The rationales are structured as a set of sub-explanations that reason about documents’ shared authorship in terms of Av-relevant linguistic features Hung et al. (2023); Boenninghoff et al. (2019); these sub-explanations have corresponding (intermediate) labels that provide further structure and also serve as a means to verify overall consistency with Cave’s authorship prediction.
@@
Since there is no human-written Av explanation data that can be used to train our in-house model, we adopt the model distillation strategy: West et al. (2022); Li et al. (2023a, 2024) we engineer a prompt to generate rationales in the structure described above, generate silver-standard training data by prompting a large teacher model (GPT-4-Turbo OpenAI (2024a)), and use that to finetune a smaller student model (LLaMa-3-8B Meta-AI (2024)). We also introduce our automatic metric Consistency that measures the consistency between the rationale and the model’s final label; Consistency is used to filter the silver train data before finetuning Ramnath et al. (2023); Li et al. (2023a), and also assess Cave during test time. Our experiments with three difficult Av datasets show that Cave leads to competitive task performances as well as high quality of rationales.

### v3

@@ Abstract
Authorship Verification (Av) (do two documents have the same author?) is essential in many real-life applications. Av is often used in privacy-sensitive domains that require an offline proprietary model that is deployed on premises, making publicly served online models (APIs) a suboptimal choice. Current offline AV models however have lower downstream utility due to limited accuracy (eg: traditional stylometry Av systems) and lack of accessible post-hoc explanations. In this work, we address the above challenges by developing a trained, offline model Cave (Controllable Authorship Verification Explanations): Cave generates free-text Av explanations that are controlled to be (1) accessible (uniform structure that can be decomposed into sub-explanations grounded to relevant linguistic features), and (2) easily verified for explanation-label consistency. We generate silver-standard training data grounded to the desirable linguistic features by a prompt-based method Prompt-CAVE. We then filter the data based on rationale-label consistency using a novel metric Cons-R-L. Finally, we fine-tune a small, offline model (LLaMa-3-8B) with this data to create our model Cave. Results on three difficult Av datasets show that Cave generates high quality explanations (as measured by automatic and human evaluation) as well as competitive task accuracy111github.com/INK-USC/Controllable-AV-Explanations.
@@
Sahana $\varheartsuit\varheartsuit{}^{\varheartsuit}$ Kartik $\varheartsuit\varheartsuit{}^{\varheartsuit}$ Elizabeth Boschee Xiang $\varheartsuit\varheartsuit{}^{\varheartsuit}$$\varheartsuit{}^{\varheartsuit}$epartment of Computer Science, University of Southern California Information Sciences Institute, University of Southern California sramnath@usc.edu
@@ 1 Introduction
Authorship Verification (Av) Koppel and Winter (2014) is the NLP task of determining if two input documents were written by the same author. Av is used for tasks such as plagiarism detection, forensic analysis (often in support of law enforcement), analysis of the spread of misinformation. Given the sensitive nature of this task, it is imperative to develop Av methods that are explainable, accessible and secure. Explainability ensures that the system’s decisions can be understood and trusted by the user. Accessibility ensures easier downstream usability; for example, a well-structured free-text explanation with clearly defined topic points is more easily understood than unstructured explanations such as long lists of low-level features. Finally, security is necessary to safely use the Av system for proprietary data; a locally-hosted model on the user’s server is more secure than an API call to an online model, which could potentially store the user’s sensitive data. Av is also an extremely difficult task for humans, in contrast to other NLP tasks such as machine translation; this further necessitates the need for automatic, high-quality Av systems.
Figure 1: Cave generates uniformly structured free-text explanations grounded in relevant linguistic features, that can be automatically verified for consistency.
special characters style, capitalization style: Text 1 uses continuous capitalization for emphasis (e.g., ‘WONDERFUL’, ‘THRILLED’), which is not observed in Text 2. NO,
acronyms and abbreviations: Neither text makes significant use of acronyms or unusual abbreviations. YES,
expressions and idioms: Both texts avoid colloquial expressions and idioms, opting for a more formal format. YES,
final score: 0.375,
output: NO
Table 1: Output structure of Cave, Prompt-CAVE: The text (JSON keys) highlighted in blue depict the linguistic features used in the analysis, as well as the keys for final confidence score and predicted task output. The text underlined and highlighted in yellow depict the intermediate labels with respect to each linguistic feature, the confidence score value, and the predicted task label. We use the predicted label to measure task accuracy and all of them to measure rationale-label consistency.
@@
Early works performed Av by comparing hand-crafted features such as n-grams van Dam (2013), POS tags Moreau and Vogel (2013), LIWC features Uchendu et al. (2020); while these are explainable, they are hard to scale. Subsequent works employed neural architectures to improve scalability and accuracy, such as the usage of Siamese networks Araujo-Pino et al. (2020); Najafi and Tavan (2022); Boenninghoff et al. (2019) to assign authorship via distance between document embeddings. These methods, however, have little or no post-hoc explainability. This poses a huge problem: one cannot, for instance, make legal assertions about the authorship of a set of questionable documents solely on the basis of “a system said so” - decision makers (judges, university officials, intelligence analysts, etc.) need to know why. Recent years have seen the surge of large language models and their self-rationalization capabilities: Hung et al. (2023); Huang et al. (2024b) generate free-text rationales [^2] We use terms explanation and rationale interchangeably. for Av from models such as GPT-4 OpenAI (2024b) in a zero/few-shot manner. These methods however face issues such as unreliability of online models (varying results/details across runs, deprecation of APIs), inconsistency of rationales with predicted label (C), and high expenses.
@@
In this work, we propose our model Cave (Controllable Authorship Verification Explanations), an in-house model that generates authorship predictions along with high-quality, free-text rationales that explain the former post-hoc. The rationales are structured as a set of sub-explanations (refer Table 1) grounded to Av-relevant linguistic features Hung et al. (2023); Boenninghoff et al. (2019); these sub-explanations have corresponding (intermediate) labels that provide further structure and also serve as a means to verify overall consistency with Cave’s authorship prediction.
@@
Since there is no human-written Av explanation data that can be used to train our in-house model, we adopt the model distillation strategy: West et al. (2022); Li et al. (2023a, 2024) we engineer Prompt-CAVE to generate silver-standard train data in the format described above by prompting a large, oracle model (GPT-4-Turbo OpenAI (2024a)), and use that to supervised-finetune a smaller language model (LLaMa-3-8B Meta-AI (2024)). We also introduce our automatic metric Cons-R-L that measures the consistency between the rationale and the model’s final label; Cons-R-L is used to filter the silver train data before finetuning Ramnath et al. (2023); Li et al. (2023a), and also assess Cave during inference.
Figure 2: Pipeline to train Cave: We obtain silver train data from GPT-4-Turbo using Prompt-CAVE, filter it according to Cons-R-L and our output format. We then supervised-finetune a LLaMa-3-8B with the filtered data.
Our experiments with three difficult Av datasets show that Cave leads to competitive task performances as well as high quality of rationales.